OpenStudy (anonymous):

8 and 2/3 divided by 2 and 1/3

OpenStudy (anonymous):

\[3\frac{5}{7} \]

OpenStudy (anonymous):

Simplify (8+ 2/3)/ (2+1/3)

jhonyy9 (jhonyy9):

8 2/3 =26/3 2 1/3 =7/3 26/3/7/3 = 26/3 * 3/7 = 26/7

jhonyy9 (jhonyy9):

or 3 5/7